Calculate the volume of carbon(IV)oxide evolved at s.t.p by dissolving 20g of pure marble, CaCO3 in hydrochloric acid.
1
Expert's answer
2021-08-27T02:32:38-0400

CaCO3 = CO2

1 mol produce 1 mole

20 g / 100 g/mol = 0.2 mol CaCO3

So n (CO2) = 0.2 mol as well!

STP:

273.15 K

105 Pascals

1 mole of an ideal gas has a volume of 22.4 liters

V (CO2) = 0.2×22.4 = 4.48 L.